The lateral edge of a square prism is 30 cm long and it is inclined at an angle of 70 degrees with the base. If the volume of the prism is 1,500 cm3,

____________________ 2) What is the base edge in cm?

____________________ 3) What is the lateral area of the prism in cm2?

_____________________ 4) What is the total area of the square prism in cm2?
1
Expert's answer
2020-12-03T20:07:24-0500

1. "V=B\u00b2h"

B=Base edge

h=height


"h=\sin(70\degree)×30=28.2 cm"

"B=\sqrt{V \over h}=\sqrt{1500\over 28.2}=7.3cm"


2. Lateral area of the prism is "L.A.=4Bh=4\u00d77.3\u00d728.2=823.44 cm\u00b2"


3. Total area of = ph + 2A

p=perimeter of the base

h=height of the prism

A=Area of the base


p = 4 × 7.3 = 29.2cm

h = 28.2cm

A = 7.3 × 7.3 = 53.29cm²


T.A. = (29.2 × 28.2) + 53.29

= 823.44 + 53.29 = 876.73cm²